# Learning Mean-Field Games through Mean-Field Actor-Critic Flow

### v1

@@ B.2 Convergence of the gap for value function
$\leq C_{R_{2}}\left(|t-t^{\prime}|^{\frac{1}{2}}+|x-x^{\prime}|\right)$
where we used (B.75) in the second inequality. Similarly,
$\left|b(t,x,\mu^{k}_{t},\alpha^{\diamond}(t,x))-b(t^{\prime},x^{\prime},\mu^{k}_{t^{\prime}},\alpha^{\diamond}(t,x))\right|$ $\leq C_{R_{2}}\left(|t-t^{\prime}|^{\frac{1}{2}}+|x-x^{\prime}|\right)$
$\left|\sigma(t,x,\mu^{k}_{t})-\sigma(t^{\prime},x^{\prime},\mu^{k}_{t^{\prime}})\right|$ $\leq C\left(|t-t^{\prime}|+|x-x^{\prime}|\right)$
hold. We also have the bound
$\left|b(t^{\prime},x^{\prime},\mu^{k}_{t^{\prime}},\alpha^{\diamond}(t,x))\right|$ $\leq K(1+|x^{\prime}|+W_{2}(\mu^{k}_{t^{\prime}},\delta_{0})+|\alpha^{\diamond}(t,x)|)\leq C_{R_{2}},$
$\left|\sigma(t^{\prime},x^{\prime},\mu^{k}_{t^{\prime}})\right|$ $\leq K.$
By the Hölder condition (B.27), we have
$\left|\nabla_{x}V_{k}(t,x)-\nabla_{x}V_{k}(t^{\prime},x^{\prime})\right|,\,\left|\nabla_{x}^{2}V_{k}(t,x)-\nabla_{x}^{2}V_{k}(t^{\prime},x^{\prime})\right|\leq C_{R_{2}}\left(|t-t^{\prime}|^{\frac{1}{2}}+|x-x^{\prime}|\right)^{\zeta}$
$\left|\nabla_{x}V_{k}(t,x)\right|,\left|\nabla_{x}^{2}V_{k}(t,x)\right|\leq C_{R_{2}}.$
Combining all the estimations above (and $|x-x^{\prime}|,|t-t^{\prime}|\leq r_{3}<1$) into (B.78), we obtain
$\left|H(t,x,\alpha^{\diamond}(t,x))-H(t^{\prime},x^{\prime},\alpha^{\diamond}(t,x))\right|\leq C_{R_{2}}\left(|t-t^{\prime}|^{\frac{\zeta}{2}}+|x-x^{\prime}|^{\zeta}\right)\leq C_{R_{2}}r_{3}^{\frac{\zeta}{2}}.$ (B.80)
Substituting (B.80) into (B.77), we get (B.76).
Step 6. We estimate $(\text{I})$ (B.72) and $(\text{II})$ (B.73) in this step. For $(\text{II})$, we have
$(\text{II})$ $\leq K\left(1+|\widehat{x}_{k}^{\prime}|+|\alpha_{k}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime})|\vee|\alpha_{k}^{\diamond}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime})|+|\nabla_{x}V_{k}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime})|)\right)\,|\alpha_{k}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime})-\alpha_{k}^{\diamond}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime})|$ (B.81)
@@
where we used (B.74) in the first inequality, used (B.44), (B.75), and (B.27) in the second inequality, used (B.45) and (B.76) in the third inequality. We set $r_{3}$ to be small enough s.t. $C_{R_{2}}r_{3}^{\frac{\zeta}{4}}\leq\frac{\lambda}{4\Delta t^{2}}$ (cf. Figure 9), then (B.81) becomes
@@
We estimate $(\text{I})$ next. Since the local optimal $\alpha^{\diamond}_{k}$ maximize the Hamiltonian (cf. (B.3)), we can replace $\alpha_{k}^{\diamond}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime})$ by $\alpha^{*}_{k}(\widehat{s}_{k}^{\prime},\widehat{y}_{k}^{\prime})$ in $(\text{I})$ and get
@@
$\quad+\big[f(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime},\mu^{k}_{\widehat{t}_{k}^{\prime}},\alpha_{k}^{*}(\widehat{s}_{k}^{\prime},\widehat{y}_{k}^{\prime}))-f(\widehat{s}_{k}^{\prime},\widehat{y}_{k}^{\prime},\mu^{k}_{\widehat{s}_{k}^{\prime}},\alpha^{*}_{k}(\widehat{s}_{k}^{\prime},\widehat{y}_{k}^{\prime}))\big]$
$=:(\text{III})+(\text{IV})+(\text{V}).$
We estimate each term in (B.83) next. We remark that the estimation for (B.83) is similar to the left part of (B.78), because both are the difference of Hamiltonian with same input argument $\alpha$. We first observe that
$|\widehat{t}_{k}^{\prime}-\widehat{s}_{k}^{\prime}|\leq|\widehat{t}_{k}^{\prime}-\widehat{t}_{k}|+|\widehat{t}_{k}-t_{k}|+|t_{k}-s_{k}|+|s_{k}-\widehat{s}_{k}|+|\widehat{s}_{k}-\widehat{s}_{k}^{\prime}|\leq 4r_{3}+C_{1}(\gamma\Delta t)^{-\frac{2}{l-2}}(\varepsilon+\delta)$ (B.84)
$|\widehat{x}_{k}^{\prime}-\widehat{s}_{k}^{\prime}|\leq|\widehat{x}_{k}^{\prime}-\widehat{x}_{k}|+|\widehat{x}_{k}-x_{k}|+|x_{k}-y_{k}|+|y_{k}-\widehat{y}_{k}|+|\widehat{y}_{k}-\widehat{y}_{k}^{\prime}|\leq 4r_{3}+C_{1}(\gamma\Delta t)^{-\frac{2}{l-2}}(\varepsilon+\delta)$
where we used (B.45), (B.40), and (B.25).
The estimation for $(\text{V})$ is similar to (B.79), except that $|\alpha_{k}^{\diamond}(\widehat{t}_{k},\widehat{x}_{k})|\leq C_{R_{2}}$ is replaced by $|\alpha_{k}(\widehat{t}_{k},\widehat{x}_{k})|\leq C(1+|\widehat{x}_{k}|)\leq CR_{2}$. Therefore,
@@
where we used (B.84) in the lase inequality. Recall we set $\varepsilon,\delta$ to be small enough s.t. $(\gamma\Delta t)^{-\frac{2}{l-2}}(\varepsilon+\delta)\leq c_{2}\eta/(2R_{2}^{2})\leq 1$ in Step 3.
@@
$(\text{VI}):=\big[b(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime},\mu^{k}_{\widehat{t}_{k}^{\prime}},\alpha_{k}^{*}(\widehat{s}_{k}^{\prime},\widehat{y}_{k}^{\prime}))-b(\widehat{s}_{k}^{\prime},\widehat{y}_{k}^{\prime},\mu^{k}_{\widehat{s}_{k}^{\prime}},\alpha^{*}_{k}(\widehat{s}_{k}^{\prime},\widehat{y}_{k}^{\prime}))\big]^{\top}\nabla_{x}V_{k}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime})$
and
@@
$\leq\Big[CR_{2}(r_{3}^{\frac{1}{2}}+(\gamma\Delta t)^{-\frac{1}{l-2}}(\varepsilon+\delta)^{\frac{1}{2}})\Big]C_{R_{2}}=C_{R_{2}}(r_{3}^{\frac{1}{2}}+(\gamma\Delta t)^{-\frac{1}{l-2}}(\varepsilon+\delta)^{\frac{1}{2}}),$
where we used (B.27) in the first inequality, and (B.84) in the second inequality. For $(\text{VII})$, we have
@@
where we used (B.47) in the first inequality, used (B.64) and (B.66) in the second inequality (note the term with $\tfrac{1}{\delta}$ cancels), used $t_{+}-\widehat{t}_{k}+\Delta t,t_{+}-\widehat{s}_{k}+\Delta t\leq 2\Delta t$ and (B.40) in the third inequality, and Cauchy’s inequality ($lab^{l-1}\leq a^{l}+(l-1)b^{l}$) in the last inequality. Combining (B.86) and (B.87), we obtain
@@
We remark that if we estimate $(\text{IV})$ with the same strategy for estimating (B.78), the term $C\gamma\Delta t\big(|\widehat{x}_{k}|_{l}^{l}+|\widehat{x}_{k}|_{l}^{l}\big)$ will depend on $R_{2}$ and we will not be able to derive a contradiction in the next step. We estimate $(\text{III})$ next.
$(\text{III})$ $=\dfrac{1}{2}\operatorname{Tr}\Bigg[\begin{pmatrix}\sigma(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime},\mu^{k}_{\widehat{t}_{k}^{\prime}})\\ \sigma(\widehat{s}_{k}^{\prime},\widehat{y}_{k}^{\prime},\mu^{k}_{\widehat{s}_{k}^{\prime}})\end{pmatrix}^{\top}\begin{bmatrix}\nabla_{x}^{2}V_{k}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime})&0\\ 0&-\nabla_{y}^{2}V^{*}(\widehat{s}_{k}^{\prime},\widehat{y}_{k}^{\prime})\end{bmatrix}\begin{pmatrix}\sigma(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime},\mu^{k}_{\widehat{t}_{k}^{\prime}})\\ \sigma(\widehat{s}_{k}^{\prime},\widehat{y}_{k}^{\prime},\mu^{k}_{\widehat{s}_{k}^{\prime}})\end{pmatrix}\Bigg$]
@@
$\leq\dfrac{1}{2}\operatorname{Tr}\left[\begin{pmatrix}\sigma(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime},\mu^{k}_{\widehat{t}_{k}^{\prime}})\\ \sigma(\widehat{s}_{k}^{\prime},\widehat{y}_{k}^{\prime},\mu^{k}_{\widehat{s}_{k}^{\prime}})\end{pmatrix}^{\top}\left(2\gamma\Delta t\,l(l-1)\begin{bmatrix}\widehat{D}^{k}_{x}&0\\ 0&\widehat{D}^{k}_{y}\end{bmatrix}+\dfrac{1}{\delta}\begin{bmatrix}I_{n}&-I_{n}\\ -I_{n}&I_{n}\end{bmatrix}+\mu I_{2n}\right)\begin{pmatrix}\sigma(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime},\mu^{k}_{\widehat{t}_{k}^{\prime}})\\ \sigma(\widehat{s}_{k}^{\prime},\widehat{y}_{k}^{\prime},\mu^{k}_{\widehat{s}_{k}^{\prime}})\end{pmatrix}\right$]
where we used (B.47) in the first inequality, and (B.67) in the second inequality. Therefore, we have
@@
$\leq C\gamma\Delta t\left(|\widehat{x}_{k}|_{l-2}^{l-2}+|\widehat{y}_{k}|_{l-2}^{l-2}\right)+\frac{C}{\delta}\left(r_{3}^{2}+(\gamma\Delta t)^{-\frac{4}{l-2}}(\varepsilon+\delta)^{2}\right)+\mu K^{2},$
where we used (B.68) in the second inequality, used (B.84) in the last inequality. Combining (B.85), (B.88), and (B.89) into (B.83), we obtain
@@
where we used $l|x|_{l-2}^{l-2}\leq 2d+(l-2)|x|_{l}^{l}$. Recall the dependence of the parameters are set in Figure 9. Following this dependence, we set $\mu$ to be small enough s.t. $\mu K^{2}\leq\frac{\lambda}{4\Delta t^{2}}$. Then we set $\gamma$ to be small enough s.t. $C_{4}\gamma\Delta t\leq\frac{\lambda}{4\Delta t^{2}}$. Then we set $\varepsilon=\delta$ to be small enough (depending on $R_{2},\gamma,\Delta t$) s.t.
@@
where $C_{R_{2}}$ corresponds to the constants in (B.90). Then, we set $r_{3}$ to be small enough s.t. $C_{R_{2}}r_{3}^{\frac{1}{2}}+Cr_{3}^{2}/\delta\leq\frac{\lambda}{4\Delta t^{2}}$. Combining these settings together into (B.90), we obtain
@@
Step 7. We use the previous estimations to derive a contradiction in this step. We combine (B.70) and (B.92) an get
@@
Let $\Delta t\leq 1/C_{4}$, the inequality simplifies to
@@
Let $r_{3}$ be small enough s.t. $2\mu r_{3}\leq\frac{\lambda}{4\Delta t^{2}}$. Additionally, since |(q,p,q^,p^)||($\widehat{q},\widehat{p}$)| satisfies (B.61), we can set first $r_{3}$ then $\iota$ to be small enough s.t.
$C_{5}R_{2}(|p|+|\widehat{p}|)+|q|+|\widehat{q}|\leq\frac{\lambda}{4\Delta t^{2}}.$
Then, we obtain
@@
Next, we square both sides of (B.93) and integrate $(\widehat{t}_{k},\widehat{x}_{k})$ against the density function $\rho^{k}=\rho^{\mu^{\tau_{k}},\alpha^{\tau_{k}}}$ on a small domain $I_{k}\times B(\tfrac{1}{2}r_{3},x_{k})$, where $I_{k}:=[t_{k}-\frac{1}{2}r_{3},t_{k}+\frac{1}{2}r_{3}]$ and $B(\tfrac{1}{2}r_{3},x_{k}):=\{x\in\mathbb{R}^{d}~|~|x-x_{k}|\leq\frac{1}{2}r_{3}\}$. We obtain
$\quad\int_{I_{k}}\int_{B(\tfrac{1}{2}r_{3},x_{k})}\frac{\lambda^{2}}{16\Delta t^{4}}\rho^{k}(t,x)\,\mathrm{d}x\,\mathrm{d}t\leq C_{R_{2}}\int_{I_{k}}\int_{B(\tfrac{1}{2}r_{3},x_{k})}\left|\alpha_{k}(t,x)-\alpha_{k}^{\diamond}(t,x)\right|^{2}\rho^{k}(t,x)\,\mathrm{d}x\,\mathrm{d}t$ (B.94)
$\leq C_{R_{2}}\left\|\alpha_{k}-\alpha_{k}^{\diamond}\right\|_{k}^{2}\leq\frac{C_{R_{2}}}{k^{2}}\left\|\alpha_{k}-\alpha_{k}^{*}\right\|_{k}^{2}\leq\frac{C_{R_{2}}}{k^{2}}$
where the third inequality is because of the assumption (B.17), and the last inequality is because $\left\|\alpha_{k}-\alpha_{k}^{*}\right\|_{k}$ is bounded. Since the density function $\rho^{k}(t,x)$ has a lower bound $c_{R_{2}}$ when $|x|\leq R_{2}$, (see (4.1)), (B.94) becomes
$\frac{C_{R_{2}}}{k^{2}}\geq\int_{I_{k}}\int_{B(\tfrac{1}{2}r_{3},x_{k})}\frac{\lambda^{2}}{16\Delta t^{4}}\rho^{k}(t,x)\,\mathrm{d}x\,\mathrm{d}t\geq|I_{k}|\,|B(\tfrac{1}{2}r_{3},x_{k})|\,\frac{\lambda^{2}}{16\Delta t^{4}}\,c_{R_{2}}=Cr_{3}^{d+1}\,\frac{\lambda^{2}}{16\Delta t^{4}}\,c_{R_{2}}.$
Finally, we reach a contradiction by letting $k\to\infty$. This contradiction comes from the assumption that $\{h_{k}(t)\}_{k=1}^{\infty}$ has a subsequence that converges to a nonzero function $h(t)$. Therefore,
$\limsup_{k\to\infty}h(t)=0\quad\forall t\in[0,T],$
and hence (B.18) holds. ∎

### v2

@@ B.2 Convergence of the gap for value function
$\leq C_{R_{2}}\left(|t-t^{\prime}|^{\frac{1}{2}}+|x-x^{\prime}|\right),$
where we use (B.70) in the second inequality. Similarly,
$\left|b(t,x,\mu^{k}_{t},\alpha^{\diamond}(t,x))-b(t^{\prime},x^{\prime},\mu^{k}_{t^{\prime}},\alpha^{\diamond}(t,x))\right|$ $\leq C_{R_{2}}\left(|t-t^{\prime}|^{\frac{1}{2}}+|x-x^{\prime}|\right),$
$\left|\sigma(t,x,\mu^{k}_{t})-\sigma(t^{\prime},x^{\prime},\mu^{k}_{t^{\prime}})\right|$ $\leq C\left(|t-t^{\prime}|+|x-x^{\prime}|\right).$
Using bounds $\left|b(t^{\prime},x^{\prime},\mu^{k}_{t^{\prime}},\alpha^{\diamond}(t,x))\right|\leq C_{R_{2}}$, $\left|\sigma(t^{\prime},x^{\prime},\mu^{k}_{t^{\prime}})\right|\leq K$, $|x-x^{\prime}|\vee|t-t^{\prime}|\leq r_{3}<1$ (cf. (B.43)), the Hölder condition (B.26), and all the estimations above, (B.73) becomes
$\left|H(t,x,\alpha^{\diamond}(t,x))-H(t^{\prime},x^{\prime},\alpha^{\diamond}(t,x))\right|\leq C_{R_{2}}\left(|t-t^{\prime}|^{\frac{\zeta}{2}}+|x-x^{\prime}|^{\zeta}\right)\leq C_{R_{2}}r_{3}^{\frac{\zeta}{2}},$ (B.75)
which concludes the proof of (B.71).
Step 6. We estimate $(\text{I})$ (B.67) and $(\text{II})$ (B.68) respectively. For the term $(\text{II})$, by (B.69),
$(\text{II})$ $\leq C\left(1+|\widehat{x}_{k}^{\prime}|+|\alpha_{k}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime})|\vee|\alpha_{k}^{\diamond}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime})|+|\nabla_{x}V_{k}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime})|\right)\,|\alpha_{k}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime})-\alpha_{k}^{\diamond}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime})|$ (B.76)
@@
where the second inequality follows from (B.26), (B.42), and (B.70), while the third inequality follows from (B.43) and (B.71). Set $r_{3}$ to be small enough such that $C_{R_{2}}r_{3}^{\frac{\zeta}{4}}\leq\frac{\lambda}{4\Delta t^{2}}$ (cf. Figure 9). We get
@@
Next, we estimate $(\text{I})$. By the definition of $\alpha^{\diamond}_{k}$ (cf. (B.3)),
@@
$\quad+\big[f(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime},\mu^{k}_{\widehat{t}_{k}^{\prime}},\alpha_{k}^{*}(\widehat{s}_{k}^{\prime},\widehat{y}_{k}^{\prime}))-f(\widehat{s}_{k}^{\prime},\widehat{y}_{k}^{\prime},\mu^{k}_{\widehat{s}_{k}^{\prime}},\alpha^{*}_{k}(\widehat{s}_{k}^{\prime},\widehat{y}_{k}^{\prime}))\big]=:(\text{III})+(\text{IV})+(\text{V}).$
We estimate each term in (B.78) separately. We remark that, the estimation for (B.78) is similar to that in (B.73), both being the difference of Hamiltonian with the same input argument $\alpha$. Note that
$|\widehat{t}_{k}^{\prime}-\widehat{s}_{k}^{\prime}|\leq|\widehat{t}_{k}^{\prime}-\widehat{t}_{k}|+|\widehat{t}_{k}-t_{k}|+|t_{k}-s_{k}|+|s_{k}-\widehat{s}_{k}|+|\widehat{s}_{k}-\widehat{s}_{k}^{\prime}|\leq 4r_{3}+C_{1}(\gamma\Delta t)^{-\frac{2}{l-2}}(\varepsilon+\delta),$ (B.79)
$|\widehat{x}_{k}^{\prime}-\widehat{y}_{k}^{\prime}|\leq|\widehat{x}_{k}^{\prime}-\widehat{x}_{k}|+|\widehat{x}_{k}-x_{k}|+|x_{k}-y_{k}|+|y_{k}-\widehat{y}_{k}|+|\widehat{y}_{k}-\widehat{y}_{k}^{\prime}|\leq 4r_{3}+C_{1}(\gamma\Delta t)^{-\frac{2}{l-2}}(\varepsilon+\delta),$
which follow from (B.25), (B.38), and (B.43).
The estimation for $(\text{V})$ is similar to (B.74), except that $|\alpha_{k}^{\diamond}(\widehat{t}_{k},\widehat{x}_{k})|\leq C_{R_{2}}$ is replaced by $|\alpha_{k}(\widehat{t}_{k},\widehat{x}_{k})|\leq C(1+|\widehat{x}_{k}|)\leq CR_{2}$. By (B.79),
@@
Recall that we have set $\varepsilon,\delta$ to be small enough such that $C_{1}(\gamma\Delta t)^{-\frac{2}{l-2}}(\varepsilon+\delta)\leq c_{2}\eta/(2R_{2}^{2})$ in Step 3.
@@
$(\text{VI}):=\big[b(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime},\mu^{k}_{\widehat{t}_{k}^{\prime}},\alpha_{k}^{*}(\widehat{s}_{k}^{\prime},\widehat{y}_{k}^{\prime}))-b(\widehat{s}_{k}^{\prime},\widehat{y}_{k}^{\prime},\mu^{k}_{\widehat{s}_{k}^{\prime}},\alpha^{*}_{k}(\widehat{s}_{k}^{\prime},\widehat{y}_{k}^{\prime}))\big]^{\top}\nabla_{x}V_{k}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime}),$
@@
$\leq\Big[CR_{2}(r_{3}^{\frac{1}{2}}+(\gamma\Delta t)^{-\frac{1}{l-2}}(\varepsilon+\delta)^{\frac{1}{2}})\Big]C_{R_{2}}=C_{R_{2}}(r_{3}^{\frac{1}{2}}+(\gamma\Delta t)^{-\frac{1}{l-2}}(\varepsilon+\delta)^{\frac{1}{2}}).$
For $(\text{VII})$, by (B.44), (B.60), (B.62) and (B.38),
@@
where the last inequality follows from Young’s inequality $ab^{l-1}\leq\frac{1}{l}a^{l}+\frac{l-1}{l}b^{l},\ \forall a,b>0$. Combining (B.81) and (B.82) yields
@@
We remark that, if $(\text{IV})$ is estimated using the same strategy as that in (B.73), the constant $C$ in $C\gamma\Delta t\big(|\widehat{x}_{k}|_{l}^{l}+|\widehat{y}_{k}|_{l}^{l}\big)$ will depend on $R_{2}$, causing the failure to reaching a contradiction in the next step.
For $(\text{III})$, by (B.44) and (B.63),
$(\text{III})$ $=\dfrac{1}{2}\operatorname{Tr}\Bigg[\begin{pmatrix}\sigma(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime},\mu^{k}_{\widehat{t}_{k}^{\prime}})\\ \sigma(\widehat{s}_{k}^{\prime},\widehat{y}_{k}^{\prime},\mu^{k}_{\widehat{s}_{k}^{\prime}})\end{pmatrix}^{\top}\begin{bmatrix}\nabla_{x}^{2}V_{k}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime})&0\\ 0&-\nabla_{y}^{2}V_{k}^{*}(\widehat{s}_{k}^{\prime},\widehat{y}_{k}^{\prime})\end{bmatrix}\begin{pmatrix}\sigma(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime},\mu^{k}_{\widehat{t}_{k}^{\prime}})\\ \sigma(\widehat{s}_{k}^{\prime},\widehat{y}_{k}^{\prime},\mu^{k}_{\widehat{s}_{k}^{\prime}})\end{pmatrix}\Bigg$]
@@
$\leq\dfrac{1}{2}\operatorname{Tr}\left[\begin{pmatrix}\sigma(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime},\mu^{k}_{\widehat{t}_{k}^{\prime}})\\ \sigma(\widehat{s}_{k}^{\prime},\widehat{y}_{k}^{\prime},\mu^{k}_{\widehat{s}_{k}^{\prime}})\end{pmatrix}^{\top}\left(2\gamma\Delta t\,l(l-1)\begin{bmatrix}\widehat{D}^{k}_{x}&0\\ 0&\widehat{D}^{k}_{y}\end{bmatrix}+\dfrac{1}{\delta}\begin{bmatrix}I_{n}&-I_{n}\\ -I_{n}&I_{n}\end{bmatrix}+\mu I_{2n}\right)\begin{pmatrix}\sigma(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime},\mu^{k}_{\widehat{t}_{k}^{\prime}})\\ \sigma(\widehat{s}_{k}^{\prime},\widehat{y}_{k}^{\prime},\mu^{k}_{\widehat{s}_{k}^{\prime}})\end{pmatrix}\right$].
Therefore, by (B.64) and (B.79),
@@
$\leq C\gamma\Delta t\left(|\widehat{x}_{k}|_{l-2}^{l-2}+|\widehat{y}_{k}|_{l-2}^{l-2}\right)+\frac{C}{\delta}\left(r_{3}^{2}+(\gamma\Delta t)^{-\frac{4}{l-2}}(\varepsilon+\delta)^{2}\right)+\mu K^{2}.$
Substituting (B.80), (B.83), and (B.84) into (B.78) yields
@@
where we use $l|x|_{l-2}^{l-2}\leq 2d+(l-2)|x|_{l}^{l}$. We record the constant $C_{4}$ for parameter specification.
Recall the parameter dependence illustrated in Figure 9. Following this dependence, we set $\mu$ to be small enough such that $\mu K^{2}\leq\frac{\lambda}{4\Delta t^{2}}$. Then we set $\gamma$ to be small enough such that $C_{4}\gamma\Delta t\leq\frac{\lambda}{4\Delta t^{2}}$. Then we set $\varepsilon=\delta$ to be small enough (depending on $R_{2},\gamma,\Delta t$) such that
@@
where $C_{R_{2}}$ corresponds to the constants in (B.85). Lastly, we set $r_{3}$ to be small enough such that $C_{R_{2}}r_{3}^{\frac{1}{2}}+Cr_{3}^{2}/\delta\leq\frac{\lambda}{4\Delta t^{2}}$. Combining these settings together into (B.85), we obtain
@@
Step 7. We combine previous estimations to reach a contradiction. By (B.65) and (B.87),
@@
Setting $\Delta t\leq 1/C_{4}$, the inequality simplifies to
@@
Let $r_{3}$ be small enough such that $2\mu r_{3}\leq\frac{\lambda}{4\Delta t^{2}}$. Additionally, since |(q,p,q^,p^)||($\widehat{q},\widehat{p}$)| satisfies (B.57), we can always first find $r_{3}$, then find $\iota$ such that $C_{5}R_{2}(|p|+|\widehat{p}|)+|q|+|\widehat{q}|\leq\frac{\lambda}{4\Delta t^{2}}.$ As a result,
@@
Squaring both sides of (B.88) and integrating $(\widehat{t}_{k},\widehat{x}_{k})$ with respect to the density function $\rho^{k}=\rho^{\mu^{\tau_{k}},\alpha^{\tau_{k}}}$ on a small domain $I_{k}\times B(x_{k},\tfrac{1}{2}r_{3})$, where $I_{k}:=[t_{k}-\frac{1}{2}r_{3},t_{k}+\frac{1}{2}r_{3}]$ and $B(x_{k},\tfrac{1}{2}r_{3}):=\{x\in\mathbb{R}^{d}~|~|x-x_{k}|\leq\frac{1}{2}r_{3}\}$, yields
$\quad\int_{I_{k}}\int_{B(x_{k},\tfrac{1}{2}r_{3})}\frac{\lambda^{2}}{16\Delta t^{4}}\rho^{k}(t,x)\,\mathrm{d}x\,\mathrm{d}t\leq C_{R_{2}}\int_{I_{k}}\int_{B(x_{k},\tfrac{1}{2}r_{3})}\left|\alpha_{k}(t,x)-\alpha_{k}^{\diamond}(t,x)\right|^{2}\rho^{k}(t,x)\,\mathrm{d}x\,\mathrm{d}t$ (B.89)
$\leq C_{R_{2}}\left\|\alpha_{k}-\alpha_{k}^{\diamond}\right\|_{k}^{2}\leq\frac{C_{R_{2}}}{k^{2}}\left\|\alpha_{k}-\alpha_{k}^{*}\right\|_{k}^{2}\leq\frac{C_{R_{2}}}{k^{2}},$
where the third inequality is based on the condition (B.17), and the last inequality is based on the (uniform in $k$) boundedness of $\left\|\alpha_{k}-\alpha_{k}^{*}\right\|_{k}^{2}$, as implied by Assumption 4.3. Since the density function $\rho^{k}(t,x)$ has a lower bound $c_{R_{2}}>0$ when $|x|\leq R_{2}$, which is uniform in $k$ (see (4.1)), (B.89) becomes
$\frac{C_{R_{2}}}{k^{2}}\geq\int_{I_{k}}\int_{B(x_{k},\tfrac{1}{2}r_{3})}\frac{\lambda^{2}}{16\Delta t^{4}}\rho^{k}(t,x)\,\mathrm{d}x\,\mathrm{d}t\geq|I_{k}|\,|B(\tfrac{1}{2}r_{3},x_{k})|\,\frac{\lambda^{2}}{16\Delta t^{4}}\,c_{R_{2}}=Cr_{3}^{d+1}\,\frac{\lambda^{2}}{16\Delta t^{4}}\,c_{R_{2}}.$
Setting $k\to\infty$ provides a contradiction, which builds upon the assumption that $\{h_{k}(t)\}_{k=1}^{\infty}$ has a subsequence that converges to a nonzero function $h(t)$. Therefore, $\limsup_{k\to\infty}h_{k}(t)=0,\ \forall t\in[0,T],$ and this concludes the proof of (B.18). ∎